OpenStudy (anonymous):

Which of the following equations can be used to find a woman's present age is she is now 6 times as old as her son and next year her age will be equal to the square of her son's age? A. 6w+1=w^2+1 B. 6(w+1)=w^2+1 C. 6(w+1)=(w+1)^2 D. 6w+1=(w+1)^2 E. w+6=(w+1)^2

OpenStudy (agent0smith):

Assuming w is her son's age, which is not very smart, because woman begins with w...: 6 times as old as her son and next year her age means 6w+1 (add 1 for it being next year) the square of her son's age next year w^2 + 1

OpenStudy (anonymous):

@agent0smith next year, square of her son's age shouldn't be (w+1)^2?

OpenStudy (agent0smith):

That's what i thought i wrote, lol. It's 5:30am

OpenStudy (anonymous):

lol get some sleep man :) we need you

OpenStudy (radar):

Let w = woman's age now then son's age is w/6 now. w + 1 = woman's age next year (w/6) + 1 = son's age next year. Stated by problem: w + 1 = ((w/6) + 1)^2 w + 1 =(w^2)/36 + ((2w)/6)) + 1 Multiply thru by 36 36w + 36 = w^2 + 12w + 36 subtracting 36 from both sides 36w = w^2 + 12w Subtract 36w from both sides 0 =w^2 - 24w w(w - 24) = 0 w =24 son = 24/6 = 4

OpenStudy (radar):

Next year woman will be 25 and son will be 5

OpenStudy (radar):

5^2 = 25 25=25

OpenStudy (anonymous):

Ok, thanks for helping me out!